# Xinxu Wei1       Kanhao Zhao1       Yong Jiao1       Lifang He1       Yu Zhang [^1] footnotemark: 1

### v1

@@ 3.3 Few-Shot Graph Prompt-Tuning via Meta Learning Optimization
As illustrated in Figure 1(c), after completing the pre-training stage, we need to fine-tune the pre-trained FM to various downstream tasks, including different atlases and disorders. However, for fMRI data, traditional full-parameter fine-tuning faces two major limitations. (1). The collection of fMRI data for neurological and psychiatric disorders is often time-consuming and labor-intensive, and for some rare diseases, only a very limited number of samples are available. When performing full-parameter fine-tuning on a large-scale foundation model with limited data, the optimization of model parameters becomes insufficient, leading to significant performance degradation. (2). Full-parameter fine-tuning requires substantial training time and computational resources, making it less practical in resource-constrained environments. Therefore, we introduce graph prompts Sun et al. (2023) to prompt-tune to our BrainGFM to different diseases and atlases. Following prior work on graph prompt learning, we design brain graph prompts specifically for brain graphs, with a structure consistent with the input brain graphs. Each node in the graph prompt is a learnable parameter, and the collection of all nodes forms a learnable vector set. For the edges, we initialize a fully learnable edge matrix, where each entry is also trainable. This design allows the graph prompt to flexibly adapt the FM to various downstream tasks without modifying the backbone parameters.
@@
To optimize the parameters of our brain graph prompts, we introduce meta-learning to train the graph prompts. Specifically, we construct a multi-task dataset in which each task corresponds to a different brain disorder and atlas pair. By adopting this meta-learning paradigm, the optimized graph prompts can be flexibly transferred to unseen diseases and atlases, enabling effective adaptation using only a small number of samples from the few-shot downstream tasks. During the meta-learning optimization process, all parameters of the pre-trained model are kept frozen, and only the graph prompt parameters, which are relatively lightweight, are updated. This design enables fast tuning and adaptation. Moreover, the few-shot sample setting is particularly well-suited for optimizing the small number of graph prompt parameters; in contrast, using limited samples to fine-tune a large pre-trained foundation model with numerous parameters would lead to insufficient training and severe underfitting. The task/disorder-specific features related to each disorder, atlas, or parcellation are thus captured and stored entirely within the well-trained brain graph prompts. As a result, with the help of the learned task-specific graph prompts, the frozen BrainGFM remains consistently ready to be efficiently and rapidly adapted via prompt-tuning to unseen tasks, datasets, disorders and atlases, even when only a few samples (few-shot settings) are available. More details about the meta-learning datasets split and training procedure can be found in the Appendix I and Table 8.
@@ 3.4 Zero-Shot Graph/Language Prompt-Tuning
Building on the few-shot capability, we further introduce language prompts to enable more generalized zero-shot learning by jointly guiding both the graph prompt and the pre-trained foundation model. In zero-shot scenarios, the parameters of the graph prompt are also frozen, meaning the model cannot rely on prompt adaptation through learning. Instead, the language prompt provides semantic guidance, allowing the model to generalize and adapt to unseen downstream data, tasks, and disorder types without any gradient-based updates. As shown in Figure 1(d), in order to enable the model to recognize and distinguish between different tasks and disorders in zero-shot settings, we introduce Task/Disorder Tokens, denoted as [T/D], during the downstream fine-tuning stage, following a similar design with Atlas/Parcellation Tokens [A/P]. To construct the [T/D] tokens, we first generate a textual description for each disorder, including its full name, abbreviation, and a concise clinical summary. For example, for Major Depressive Disorder, the corresponding text description is: Major Depressive Disorder (MDD) is a common mental illness characterized by persistent and profound low mood, loss of interest, and cognitive impairment, significantly affecting daily life and social functioning.” Otte et al. (2016) We then encode these textual descriptions using a BERT model Devlin et al. (2019) pre-trained on large-scale medical corpora, such as BioClinicalBERT Huang et al. (2019); Alsentzer et al. (2019), to obtain semantic-rich text embeddings. These embeddings are subsequently projected and embedded as [T/D] tokens, which are incorporated into the model during downstream adaptation. Similarly, the construction of [A/P] tokens is also based on language text. For each atlas and parcellation, we provide a textual description of its name, such as “Schaefer100”, “Schaefer200”, or “AAL116”. These text descriptions are then encoded using the BioClinicalBERT pre-trained model to extract language embeddings, which are subsequently transformed into [A/P] tokens. As shown in Figure 1(d), the [T/D] and [A/P] tokens are ultimately concatenated with the ROI tokens from the graph embeddings as language prompt tokens. This combined input is then fed into the foundation model to guide feature extraction specific to the given dataset, task, and disorder. By introducing disorder-specific semantic priors through the [T/D] and [A/P] tokens, the model is better equipped to capture characteristics from different tasks, disorders, atlas and parcellations, thereby improving its downstream adaption ability in zero-shot settings without any training.
@@ 4.1 Comparison with Other Methods
Datasets. To demonstrate the superiority of our BrainGFM, we conducted comparative experiments. Specifically, 10 common types of neurological and psychiatric disorders were selected from a total of 25 disorders, spanning 6 datasets among the 27 datasets we collected. More details about benchmarks and datasets can be found in Appendix N. Baselines. We compare our method against a series of baseline models. Based on the data representation type, these baselines are categorized into three groups: time-series-based methods, ROI-based methods, and graph-based methods. Based on the training paradigm, they are divided into two groups: non-pre-trained FMs and pre-trained FMs. All pre-trained models are retrained on our collected pre-training dataset to ensure a fair comparison. More details about baselines can be found in Appendix M. Metrics. We evaluate all methods using four metrics: AUC, accuracy (ACC), sensitivity (SEN), and specificity (SPE). More detailed information on disorders, datasets, and benchmarks can be found in the supplementary material. As shown in Table 1, our method outperforms all previous approaches and achieves state-of-the-art performance. The pre-trained FMs significantly outperforms models without pre-training. Our method, built upon a graph transformer backbone, substantially surpasses ROI-based brain FMs (BrainMass and BrainNPT), and also outperforms time-series brain FMs (BrainLM).
@@
Method Pre-Trained ADHD200 (ADHD) ABIDE II (ASD) ADNI 2 (AD) HBN (MDD) HBN (ANX)
@@
Vanilla GCN 61.6 63.7 59.0 63.3 63.6 65.6 60.7 63.3 68.4 70.2 64.3 72.2 70.2 71.8 68.9 71.7 75.6 78.2 79.6 72.4
BrainGNN 60.5 63.3 63.1 57.3 61.2 62.7 57.3 65.2 69.1 70.6 66.9 73.4 68.8 73.0 71.4 65.7 76.3 77.2 72.7 80.5
Vanilla TF 62.4 64.6 60.8 62.1 65.3 65.2 64.2 61.9 71.7 74.4 68.6 74.2 74.1 76.6 71.3 78.7 77.8 80.9 81.6 73.9
Graph TF 64.6 65.3 62.1 66.3 65.2 67.1 67.7 63.6 73.3 76.6 70.4 75.9 76.4 77.8 73.1 80.2 80.5 81.8 77.9 83.5
BrainNetTF 63.3 64.6 65.2 61.1 66.5 66.6 66.9 65.7 74.3 76.4 77.2 70.7 75.7 74.9 72.6 78.5 78.4 81.1 75.5 80.9
BrainNPT 62.3 66.5 61.6 59.2 65.5 67.3 65.8 71.6 70.6 75.7 65.2 75.3 72.8 74.1 68.5 76.7 75.6 77.2 79.1 71.4
BrainLM 66.3 68.6 62.7 70.4 68.5 70.2 65.6 71.1 76.7 81.5 72.4 80.5 75.1 81.9 81.9 69.4 82.3 83.6 77.6 87.2
BrainMass 65.5 66.1 63.6 69.9 67.3 68.5 69.3 64.8 76.6 79.3 71.5 80.1 75.7 78.5 82.2 74.4 79.8 82.7 79.0 80.6
BrainGFM (Ours) 70.6 72.2 67.3 73.4 71.2 73.5 70.4 69.8 80.3 85.1 76.2 84.4 83.6 85.5 85.8 77.9 85.2 86.3 87.7 82.6
Method Pre-Trained HBN (OCD) HBN (PTSD) SubMex_CUD (CUD) UCLA_CNP (SCHZ) UCLA_CNP (BP)
@@
Vanilla GCN 69.7 76.9 61.3 76.6 77.5 78.7 82.1 71.8 63.8 66.2 55.6 72.5 77.2 79.5 72.3 81.8 64.3 69.6 70.1 58.2
BrainGNN 70.1 74.6 76.7 66.4 75.2 78.0 69.3 81.9 62.5 64.8 67.1 57.5 78.1 77.5 83.7 72.6 63.8 67.3 58.5 69.2
Vanilla TF 72.5 78.8 64.5 78.2 78.2 80.1 71.7 85.9 65.6 66.7 60.4 71.1 76.8 80.2 71.9 80.4 65.1 70.6 69.4 61.7
Graph TF 73.6 81.5 66.7 77.5 77.1 81.6 74.3 80.2 67.5 69.8 59.4 75.7 78.9 80.8 75.1 81.6 67.4 70.2 70.5 64.3
BrainNetTF 74.7 80.4 79.2 69.1 79.4 81.9 74.6 84.6 66.8 67.5 61.7 71.4 75.1 77.3 73.5 77.9 68.5 71.6 73.4 64.8
BrainNPT 71.2 75.3 63.6 78.4 76.4 80.6 70.3 82.5 63.7 64.2 57.5 69.3 75.0 76.5 68.7 81.9 65.5 68.7 60.2 70.1
BrainLM 78.2 83.3 72.6 84.5 80.2 84.8 81.5 73.6 68.9 71.3 63.2 73.1 81.6 82.1 77.4 85.5 70.3 73.5 74.7 66.3
BrainMass 76.8 81.2 80.6 72.5 78.1 82.5 73.2 82.4 67.5 69.4 71.7 64.0 80.8 81.2 76.5 84.2 69.7 72.3 65.8 73.3
BrainGFM (Ours) 80.4 85.8 86.7 78.5 83.2 86.3 79.5 87.4 71.1 74.6 67.7 75.5 84.2 86.7 80.4 87.9 73.5 76.3 69.6 78.2
@@ 4.4 Comparison Among Time-Series, ROI and Graph-Based Foundation Models
We compare four types of brain FMs: time-series-based FM (e.g., BrainLM), ROI-based FM (e.g., BrainMass), vanilla graph-based FM, and our proposed graph-based model, BrainGFM. The comparison spans five key dimensions: model performance, pre-training and fine-tuning efficiency, memory usage, and model complexity.
@@
In terms of performance, BrainLM achieves the best results on AUC and ACC due to its direct modeling of raw fMRI time series, effectively capturing both temporal and spatial patterns. BrainMass, which relies on static ROI features without modeling inter-regional interactions, performs the worst. The vanilla graph-based model shows intermediate performance by explicitly modeling ROI connectivity. BrainGFM, which incorporates fMRI-specific enhancements such as graph prompts and structural encodings, significantly outperforms the vanilla graph model and matches or exceeds the performance of time-series-based models. For computational efficiency, ROI-based models are the fastest in both pre-training and fine-tuning, given their compact input and lack of spatiotemporal modeling. Time-series models are the slowest due to the cost of processing long, high-dimensional sequences. Graph-based models, including BrainGFM, lie in between. Notably, BrainGFM achieves fast fine-tuning via prompt tuning while maintaining pre-training efficiency similar to the vanilla graph FM, surpassing even ROI-based models in fine-tuning speed. Regarding resource consumption, time-series models are the most memory- and compute-intensive. ROI-based models are the most lightweight. Graph-based models, while slightly more demanding than ROI-based ones due to edge computations, remain significantly more efficient than time-series models. Overall, this evaluation highlights the trade-offs between different brain modeling paradigms and how input representations, time series, ROI features, or graphs, affect both the effectiveness and efficiency of large-scale brain FMs.

### v2

@@ 3.3 Few-Shot Graph Prompt-Tuning via Meta Learning Optimization
As illustrated in Figure 1(c), after completing the pre-training stage, we need to fine-tune the pre-trained FM to various downstream tasks, including different atlases and disorders. However, for fMRI data, traditional full-parameter fine-tuning faces two major limitations. (1). The collection of fMRI data for neurological and psychiatric disorders is often time-consuming and labor-intensive, and for some rare diseases, only a very limited number of samples are available. When performing full-parameter fine-tuning on a large-scale foundation model with limited data, the optimization of model parameters becomes insufficient, leading to significant performance degradation. (2). Full-parameter fine-tuning requires substantial training time and computational resources, making it less practical in resource-constrained environments. Therefore, we introduce graph prompts (Sun et al., 2023) to prompt-tune to our BrainGFM to different diseases and atlases. Following prior work on graph prompt learning, we design brain graph prompts specifically for brain graphs, with a structure consistent with the input brain graphs. Each node in the graph prompt is a learnable parameter, and the collection of all nodes forms a learnable vector set. For the edges, we initialize a fully learnable edge matrix, where each entry is also trainable. This design allows the graph prompt to flexibly adapt the FM to various downstream tasks without modifying the backbone parameters.
@@
To optimize the parameters of our brain graph prompts, we introduce meta-learning to train the graph prompts. Specifically, we construct a multi-task dataset in which each task corresponds to a different brain disorder and atlas pair. By adopting this meta-learning paradigm, the optimized graph prompts can be flexibly transferred to unseen diseases and atlases, enabling effective adaptation using only a small number of samples from the few-shot downstream tasks. During the meta-learning optimization process, all parameters of the pre-trained model are kept frozen, and only the graph prompt parameters, which are relatively lightweight, are updated. This design enables fast tuning and adaptation. Moreover, the few-shot sample setting is particularly well-suited for optimizing the small number of graph prompt parameters; in contrast, using limited samples to fine-tune a large pre-trained foundation model with numerous parameters would lead to insufficient training and severe overfitting. The task/disorder-specific features related to each disorder, atlas, or parcellation are thus captured and stored entirely within the well-trained brain graph prompts. As a result, with the help of the learned task-specific graph prompts, the frozen BrainGFM remains consistently ready to be efficiently and rapidly adapted via prompt-tuning to unseen tasks, datasets, disorders and atlases, even when only a few samples (few-shot settings) are available. More details about the meta-learning datasets split and training procedure can be found in the Appendix L and Table 12.
@@ 3.4 Zero-Shot Graph/Language Prompt-Tuning
Building on the few-shot capability, we further introduce language prompts to enable more generalized zero-shot learning by jointly guiding both the graph prompt and the pre-trained foundation model. In zero-shot scenarios, the parameters of the graph prompt are also frozen, meaning the model cannot rely on prompt adaptation through learning. Instead, the language prompt provides semantic guidance, allowing the model to generalize and adapt to unseen downstream data, tasks, and disorder types without any gradient-based updates. As shown in Figure 1(d), in order to enable the model to recognize and distinguish between different tasks and disorders in zero-shot settings, we introduce Task/Disorder Tokens, denoted as [T/D], during the downstream fine-tuning stage, following a similar design with Atlas/Parcellation Tokens [A/P]. To construct the [T/D] tokens, we first generate a textual description for each disorder, including its full name, abbreviation, and a concise clinical summary. For example, for Major Depressive Disorder, the corresponding text description is: Major Depressive Disorder (MDD) is a common mental illness characterized by persistent and profound low mood, loss of interest, and cognitive impairment, significantly affecting daily life and social functioning.” (Otte et al., 2016) We then encode these textual descriptions using a BERT model (Devlin et al., 2019) pre-trained on large-scale medical corpora, such as BioClinicalBERT (Huang et al., 2019; Alsentzer et al., 2019), to obtain semantic-rich text embeddings. These embeddings are subsequently projected and embedded as [T/D] tokens, which are incorporated into the model during downstream adaptation. Similarly, the construction of [A/P] tokens is also based on language text. For each atlas and parcellation, we provide a textual description of its name, such as “Schaefer100”, “Schaefer200”, or “AAL116”. These text descriptions are then encoded using the BioClinicalBERT pre-trained model to extract language embeddings, which are subsequently transformed into [A/P] tokens. As shown in Figure 1(d), the [T/D] and [A/P] tokens are ultimately concatenated with the ROI tokens from the graph embeddings as language prompt tokens. This combined input is then fed into the foundation model to guide feature extraction specific to the given dataset, task, and disorder. By introducing disorder-specific semantic priors through the [T/D] and [A/P] tokens, the model is better equipped to capture characteristics from different tasks, disorders, atlas and parcellations, thereby improving its downstream adaptation ability in zero-shot settings without any training.
@@ 4.1 Comparison with Other Methods
Datasets. To demonstrate the superiority of our BrainGFM, we conducted comparative experiments. Specifically, 10 common types of neurological and psychiatric disorders were selected from a total of 25 disorders, spanning 6 datasets among the 27 datasets we collected. More details about benchmarks and datasets can be found in Appendix R. Baselines. We compare our method against a series of baseline models. Based on the data representation type, these baselines are categorized into three groups: time-series-based methods, Connectome/FC-based methods, and graph-based methods. Based on the training paradigm, they are divided into two groups: non-pre-trained FMs and pre-trained FMs. All pre-trained models are retrained on our collected pre-training dataset to ensure a fair comparison. More details about baselines can be found in Appendix Q. Metrics. We evaluate all methods using four metrics: AUC, accuracy (ACC), sensitivity (SEN), and specificity (SPE). More detailed information on disorders, datasets, and benchmarks can be found in the supplementary material. As shown in Table 1, our method outperform all previous approaches and achieves state-of-the-art performance. The pre-trained FMs significantly outperforms models without pre-training. Our method, built upon a graph transformer backbone, substantially surpasses Connectome/FC-based brain FMs (BrainMass and BrainNPT), and also outperforms time-series brain FMs (BrainLM).
@@
Method PT ADHD200 (ADHD) ABIDE II (ASD) ADNI 2 (AD) HBN (MDD) HBN (ANX)
@@
Vanilla GCN 62.3±2.3 65.4±1.8 61.8±2.7 64.4±3.5 64.2±2.2 67.5±3.6 62.9±1.9 65.0±2.4 69.1±2.5 71.4±3.0 65.2±1.7 73.3±2.1 72.0±1.9 73.8±2.4 69.3±3.2 73.6±1.5 76.7±2.8 79.8±3.6 80.3±2.9 73.6±3.1
BrainGNN 61.3±3.1 64.4±2.6 64.0±2.9 59.9±1.8 62.7±3.5 64.9±2.1 60.2±2.4 66.0±3.2 71.1±1.6 71.2±2.9 67.6±2.5 75.3±2.8 69.0±3.1 74.1±2.2 73.9±3.0 68.5±2.5 78.7±3.3 77.0±1.8 74.2±2.9 81.8±1.6
Vanilla TF 63.7±2.8 65.9±2.4 61.9±3.3 65.6±1.9 66.4±2.1 66.5±3.2 65.7±2.6 63.4±1.8 72.9±3.0 75.8±2.3 70.2±3.5 75.4±2.6 75.5±2.0 77.4±2.7 72.7±2.9 79.3±2.4 79.5±3.8 82.9±3.3 83.4±3.1 75.3±3.0
Graph TF 64.7±3.3 66.9±2.5 63.4±2.7 67.8±2.9 66.8±1.7 68.6±2.4 69.0±3.6 64.8±2.1 74.8±2.0 77.9±3.1 71.6±3.4 77.1±2.7 75.3±2.2 78.9±2.5 74.5±3.3 76.9±2.9 81.4±3.7 83.3±2.4 79.6±3.2 85.0±2.8
BrainNetTF 64.7±2.3 66.0±3.2 66.4±2.1 62.6±3.0 67.9±1.9 68.1±2.7 68.6±2.5 67.2±3.4 75.9±2.6 77.6±2.0 78.8±2.9 72.3±3.3 74.2±2.4 76.3±2.1 73.9±3.5 79.4±2.7 80.1±3.6 82.5±2.8 76.9±3.1 82.6±2.2
BrainNPT 65.6±2.3 67.9±1.8 62.7±2.4 65.9±1.9 66.8±2.1 68.5±1.5 64.0±2.8 70.0±2.2 72.0±1.9 77.2±2.6 66.4±2.3 76.4±2.4 74.0±2.0 75.6±1.7 69.7±2.6 77.4±2.3 79.6±1.6 83.2±2.5 77.8±2.7 82.2±3.2
BrainLM 67.6±1.9 69.1±2.2 64.0±2.7 71.7±2.0 68.1±2.6 71.6±1.8 66.9±2.5 69.6±2.4 78.3±1.7 82.9±2.5 73.1±2.6 81.4±2.7 79.3±2.0 83.2±2.1 83.5±2.8 75.8±2.4 82.6±2.5 85.0±2.6 78.6±2.0 86.5±2.7
BrainMass 67.0±2.3 67.5±2.0 64.7±2.8 71.1±1.8 68.9±2.1 70.1±1.9 69.5±2.4 66.3±2.6 77.8±2.3 82.7±2.7 72.6±2.9 81.4±2.5 76.9±2.4 80.2±2.8 83.6±2.5 76.1±2.6 81.0±2.7 84.0±2.2 80.3±2.3 81.9±2.9
Brain-JEPA 69.8±1.9 71.6±2.0 66.2±1.6 72.9±2.4 70.1±1.7 73.8±1.8 70.1±1.9 69.3±1.7 79.1±2.2 84.3±1.8 76.8±2.2 83.6±2.1 83.4±1.7 85.9±1.3 84.3±1.9 77.5±1.5 85.4±1.8 86.7±2.0 85.3±2.2 83.7±1.9
BrainGFM 70.6±1.6 72.2±1.5 67.3±1.7 73.4±2.1 71.2±1.9 73.5±1.4 70.4±1.5 69.8±1.7 80.3±2.6 85.1±2.2 76.2±1.5 84.4±1.7 83.6±1.6 85.5±1.7 85.8±1.6 77.9±1.9 85.2±1.6 86.3±2.1 87.7±1.9 82.6±1.7
Method PT HBN (OCD) HBN (PTSD) SubMex_CUD (CUD) UCLA_CNP (SCHZ) UCLA_CNP (BP)
@@
Vanilla GCN 71.2±3.2 78.3±2.9 62.5±3.6 78.1±2.7 78.7±2.4 80.2±3.8 83.3±2.2 73.2±3.1 65.3±3.5 67.8±2.6 57.3±3.9 73.9±2.5 78.5±2.3 81.0±3.3 73.4±2.9 83.1±3.0 66.0±3.4 71.3±2.8 71.8±2.1 59.7±3.5
BrainGNN 71.4±2.5 76.0±3.3 78.2±2.2 67.8±3.6 76.9±3.0 79.6±3.7 70.5±2.6 83.3±3.1 63.9±3.4 66.1±3.2 68.7±2.3 59.1±3.5 79.7±2.8 78.9±2.7 85.0±3.9 74.0±2.5 65.2±3.6 68.9±3.2 60.3±2.9 70.6±3.0
Vanilla TF 73.9±2.1 80.0±2.5 65.8±3.7 79.7±3.0 77.6±3.3 81.7±3.1 73.2±2.6 87.1±2.9 67.0±3.4 68.0±2.8 61.8±3.5 72.3±3.6 78.4±2.9 81.6±3.7 73.1±2.5 82.0±3.3 66.5±3.6 72.4±2.7 70.6±3.2 63.0±2.6
Graph TF 75.0±3.1 82.8±2.6 68.0±2.9 79.1±3.2 78.4±2.3 83.1±2.7 75.5±2.1 81.7±3.1 68.8±3.2 69.4±3.3 65.0±2.8 70.0±3.6 80.1±3.0 82.2±2.9 76.8±3.3 83.2±2.7 68.7±3.5 72.3±2.4 72.0±2.9 65.6±3.4
BrainNetTF 76.1±3.4 81.9±2.7 80.4±3.1 70.7±2.8 78.8±3.0 83.2±2.5 76.2±3.2 85.9±3.1 68.3±2.9 69.1±3.4 62.9±2.6 73.0±3.7 76.4±3.1 78.7±3.0 75.0±2.5 79.3±3.3 68.1±3.1 73.2±3.4 72.6±3.6 66.3±2.8
BrainNPT 77.5±2.2 76.7±2.0 74.8±1.9 79.6±2.7 77.9±2.4 82.1±2.6 71.5±1.8 83.8±2.3 68.9±1.7 65.6±2.1 68.7±2.2 70.7±2.8 76.4±2.1 77.9±2.7 70.1±2.5 83.2±2.9 68.6±1.9 70.3±2.6 66.4±2.0 71.7±2.5
BrainLM 79.4±2.1 84.6±2.3 73.9±2.0 85.9±2.5 80.5±2.7 82.3±2.1 82.9±2.6 77.0±2.2 68.2±2.3 72.5±2.1 64.4±2.0 74.4±2.3 82.7±2.5 83.5±2.6 78.8±2.4 86.7±2.1 71.5±2.2 74.9±2.3 75.8±2.0 67.8±2.2
BrainMass 78.1±2.3 82.6±2.1 81.8±2.4 73.6±2.2 79.6±2.3 83.9±2.6 74.9±2.5 83.9±2.4 68.0±2.0 70.9±2.3 71.0±2.2 65.5±2.3 82.0±2.1 82.8±2.0 78.0±2.6 85.4±2.3 70.8±2.2 73.6±2.3 67.3±2.1 74.8±2.5
Brain-JEPA 79.3±2.4 84.1±1.5 85.2±2.2 77.4±2.1 82.2±2.4 85.7±2.3 78.5±1.6 86.9±2.1 70.1±2.0 74.9±2.3 65.9±2.2 74.1±2.1 85.0±2.5 86.6±2.4 79.6±2.2 87.3±1.3 73.7±2.2 75.2±2.1 68.8±2.4 77.8±2.5
BrainGFM 80.4±1.7 85.8±1.9 86.7±2.1 78.5±1.6 83.2±1.8 86.3±1.9 79.5±1.7 87.4±2.0 71.1±1.6 74.6±1.8 67.7±1.9 75.5±2.0 84.2±1.7 86.7±1.6 80.4±2.1 87.9±1.9 73.5±1.8 76.3±1.9 69.6±1.7 78.2±2.0
@@ 4.4 Comparison Among Time-Series, ROI and Graph-Based Foundation Models
As shown in Figure 4, we compare four types of brain FMs: time-series-based FM (e.g., BrainLM), Connectome/FC-based FM (e.g., BrainMass), vanilla graph-based FM, and our proposed graph-based model, BrainGFM. The comparison spans five key dimensions: model performance, pre-training and fine-tuning efficiency, memory usage, and model complexity.
@@
In terms of performance, BrainLM achieves the best results on AUC and ACC due to its direct modeling of raw fMRI time series, effectively capturing both temporal and spatial patterns. BrainMass, which relies on static ROI features without modeling inter-regional interactions, performs the worst. The vanilla graph-based model shows intermediate performance by explicitly modeling ROI connectivity. BrainGFM, which incorporates fMRI-specific enhancements such as graph prompts and structural encodings, significantly outperforms the vanilla graph model and matches or exceeds the performance of time-series-based models. For computational efficiency, Connectome/FC-based models are the fastest in both pre-training and fine-tuning, given their compact input and lack of spatiotemporal modeling. Time-series models are the slowest due to the cost of processing long, high-dimensional sequences. Graph-based models, including BrainGFM, lie in between. Notably, BrainGFM achieves fast fine-tuning via prompt tuning while maintaining pre-training efficiency similar to the vanilla graph FM, surpassing even Connectome/FC-based models in fine-tuning speed. Regarding resource consumption, time-series models are the most memory- and compute-intensive. Connectome/FC-based models are the most lightweight. Graph-based models, while slightly more demanding than Connectome/FC-based ones due to edge computations, remain significantly more efficient than time-series models. Overall, this evaluation highlights the trade-offs between different brain modeling paradigms and how input representations, time series, ROI features, or graphs, affect both the effectiveness and efficiency of large-scale brain FMs.